# On Weak Compactness and Uniform Regularity in Lipschitz free Spaces

### v1

@@ 5. Sequences of sums of “small molecules” contain an $\ell_{1}$-basis
Let $m\in\mathbb{N}$ be such that $\frac{1}{m}<\varepsilon$ and let $\delta>0$ be such that $\delta\left(1+\frac{1}{\varepsilon}\right)^{m}+\delta<\varepsilon D$. Next we pick $n\in\mathbb{N}$ large enough so that
@@
$\sum_{k\in I\cup J}\lambda_{k}^{n}=1-\sum_{k\notin I\cup J}\lambda_{k}^{n}=1-\sum_{k\,:\,x_{k}^{n}\in B_{i}}\lambda_{k}^{n}>1-\varepsilon.$
@@
$\left\|{\gamma+\mu}\right\|\geq\frac{D-2\varepsilon D}{D+2\varepsilon D}\left(\|\gamma\|+\left\|{\mu}\right\|\right)=\frac{1-2\varepsilon}{1+2\varepsilon}\left(\|\gamma\|+\sum_{k\in J}\lambda_{k}\right).$
@@
$=\frac{1-2\varepsilon}{1-\varepsilon}\Big(\|\gamma+\mu\|+\sum_{k\in I}\lambda_{k}^{n}\Big).$
@@
$\geq\frac{1-2\varepsilon}{1-\varepsilon}\Big(\|\gamma+\mu\|+\sum_{k\in I}\lambda_{k}^{n}\Big)-\varepsilon$
$\geq\frac{(1-2\varepsilon)^{2}}{(1-\varepsilon)(1+2\varepsilon)}\Big(\|\gamma\|+\sum_{k\in J}\lambda_{k}^{n}+\sum_{k\in I}\lambda_{k}^{n}\Big)-\varepsilon$
@@
Let $M$ be a complete pointed metric space. For every $n\in\mathbb{N}$, let $\gamma_{n}=\sum_{k}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}\in S_{\mathcal{F}(M)}$ be convex sum/series of molecules such that
$\sum_{k}d(x_{k}^{n},y_{k}^{n})\underset{n\to\infty}{\longrightarrow}0.$
@@
We assume that $W$ is bounded, otherwise it is not weakly precompact and we are done. Let $R=\sup_{\gamma\in W}\left\|{\gamma}\right\|$. Let $\mu_{n}\in\mathrm{CSM}_{\frac{1}{n}}\left({M}\right)$ be such that $\left\|{\gamma_{n}-\mu_{n}}\right\|<\frac{\left\|{\gamma_{n}}\right\|}{2}-\varepsilon$. We show that $(\gamma_{n})$ has a subsequence equivalent to the $\ell_{1}$ basis and thus $W$ is not weakly precompact. First observe that $2\varepsilon<\left\|{\gamma_{n}}\right\|\leqslant R$, so we may instead find a subsequence of $(\frac{\gamma_{n}}{\left\|{\gamma_{n}}\right\|})$ equivalent to the $\ell_{1}$-basis. Since $2\varepsilon<\left\|{\mu_{n}}\right\|$ for every $n$, Corollary 5.4 yields a subsequence of $(\frac{\mu_{n}}{\left\|{\mu_{n}}\right\|})$ which is an asymptotically isometric $\ell_{1}$-basis. Thus if we find $D<1$ such that $\left\|{\frac{\gamma_{n}}{\left\|{\gamma_{n}}\right\|}-\frac{\mu_{n}}{\left\|{\mu_{n}}\right\|}}\right\|\leqslant D$ for every $n$, Lemma 5.7 will lead to the desired conclusion. But we have by the classical Massera-Schäffer inequality [18, Tapa 15]

### v2

@@ 5. Sequences of sums of “small molecules” contain an $\ell_{1}$-basis
Let $m\in\mathbb{N}$ be such that $\frac{1}{m}\sup_{n}\sum_{k}\left|{\lambda_{k}^{n}}\right|<\varepsilon$ and let $\delta>0$ be such that $\delta\left(1+\frac{1}{\varepsilon}\right)^{m}+\delta<\varepsilon D$. Next we pick $n\in\mathbb{N}$ large enough so that
@@
$\left\|{\sum_{k\in I\cup J}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}}\right\|\geq 1-\sum_{k\notin I\cup J}\lambda_{k}^{n}=1-\sum_{k\,:\,x_{k}^{n}\in B_{i}}\lambda_{k}^{n}>1-\varepsilon.$
@@
$\left\|{\gamma+\mu}\right\|\geq\frac{D-2\varepsilon D}{D+2\varepsilon D}\left(\|\gamma\|+\left\|{\mu}\right\|\right)=\frac{1-2\varepsilon}{1+2\varepsilon}\left(\|\gamma\|+\left\|{\mu}\right\|\right).$
@@
$=\frac{1-2\varepsilon}{1-\varepsilon}\Big(\|\gamma+\mu\|+\Big\|\sum_{k\in I}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}\Big\|\Big).$
@@
$\geq\frac{1-2\varepsilon}{1-\varepsilon}\Big(\|\gamma+\mu\|+\Big\|\sum_{k\in I}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}\Big\|\Big)-\varepsilon$
$\geq\frac{(1-2\varepsilon)^{2}}{(1-\varepsilon)(1+2\varepsilon)}\Big(\|\gamma\|+\left\|{\mu}\right\|+\Big\|\sum_{k\in I}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}\Big\|\Big)-\varepsilon$
@@
Let $M$ be a complete pointed metric space. For every $n\in\mathbb{N}$, let $\gamma_{n}=\sum_{k}\lambda_{k}^{n}m_{x_{k}^{n}y_{k}^{n}}\in S_{\mathcal{F}(M)}$ such that
$\sum_{k}d(x_{k}^{n},y_{k}^{n})\underset{n\to\infty}{\longrightarrow}0\quad\text{and}\quad\sup_{n}\sum_{k}|\lambda_{k}^{n}|<\infty.$
@@
We assume that $W$ is bounded, otherwise it is not weakly precompact and we are done. Let $R=\sup_{\gamma\in W}\left\|{\gamma}\right\|$. Let $\mu_{n}\in\mathrm{CSM}_{\frac{1}{n}}\left({M}\right)$ be such that $\left\|{\gamma_{n}-\mu_{n}}\right\|<\frac{\left\|{\gamma_{n}}\right\|}{2}-\varepsilon$. We show that $(\gamma_{n})$ has a subsequence equivalent to the $\ell_{1}$ basis and thus $W$ is not weakly precompact. First observe that $2\varepsilon<\left\|{\gamma_{n}}\right\|\leqslant R$, so we may instead find a subsequence of $(\frac{\gamma_{n}}{\left\|{\gamma_{n}}\right\|})$ equivalent to the $\ell_{1}$-basis. Since $2\varepsilon<\left\|{\mu_{n}}\right\|$ for every $n$, Corollary 5.9 yields a subsequence of $(\frac{\mu_{n}}{\left\|{\mu_{n}}\right\|})$ which is an asymptotically isometric $\ell_{1}$-basis. Thus if we find $D<1$ such that $\left\|{\frac{\gamma_{n}}{\left\|{\gamma_{n}}\right\|}-\frac{\mu_{n}}{\left\|{\mu_{n}}\right\|}}\right\|\leqslant D$ for every $n$, Lemma 5.12 will lead to the desired conclusion. But we have by the classical Massera-Schäffer inequality [19, Tapa 15]